OpenStudy (lovelyanna):

Which inequality does the graph represent? A. y ≤ 1 – 3x B. y < 1 – 3x C. y > 1 – 3x D. y ≥ 1 – 3x

OpenStudy (lovelyanna):

@jigglypuff314

OpenStudy (jojo4eva):

THe answer is A it is less than and equal to because the line is solid

OpenStudy (lovelyanna):

Ok, thank you :)

OpenStudy (lovelyanna):

Is A right?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (lovelyanna):

Cool, thanks! :) can you help with some more?

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (lovelyanna):

Thanks :) Which inequality does this graph represent? A. y ≥ –2x – 4 B. y < –4x + 25 C. y ≤ – 4x + 25 D. y ≤ –2x – 4 http://static.k12.com/calms_media/media/243000_243500/243015/1/33e4ec01f9121f12c679f4a97d75b2a8442b4e71/VHS_ALG_S1_07_L314_LQ2.gif

jigglypuff314 (jigglypuff314):

solid and under therefore "y <_"

jigglypuff314 (jigglypuff314):

y-intercept = 25

OpenStudy (lovelyanna):

C?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (lovelyanna):

Thanks :)

OpenStudy (lovelyanna):

Which inequality does this graph represent? \[A: y \ge -1/2x - 2 B: y \ge 1/2x - 1 C: y \le -1/2x - 1 D: y \ge -1/2x - 1\] http://static.k12.com/calms_media/media/1312500_1313000/1312941/1/5139a5ada20768822e5fe04f48b60e3f61d25108/HS_A1G2_S1_07_14_114285.jpg

OpenStudy (lovelyanna):

Hmm.. they are all stuck together.. :/ I tried putting them apart..

OpenStudy (lovelyanna):

There are A, B, C, D, but they are stuck together, can you see them like that and read them?

jigglypuff314 (jigglypuff314):

solid and under -> "y <_" y-intercept = -1 slope = -1/2

OpenStudy (lovelyanna):

-12x - 1. the - is not negative, its subtraction, you got confused like I did, or am I misunderstanding?

OpenStudy (lovelyanna):

-1/2x - 1*

jigglypuff314 (jigglypuff314):

subtraction is adding a negtive, makes no difference "-" = + - like 1 - 2 = 1 + (-2)

OpenStudy (lovelyanna):

Oh, ok :) I know that :P Will it be C?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (lovelyanna):

Hooray! =D

OpenStudy (lovelyanna):

Which inequality does this graph represent? A. x ≥ 2 B. x ≤ 2 C. x < 2 D. x > 2 http://static.k12.com/calms_media/media/243000_243500/243017/1/97c457a46e0ae8c2fa33f18bbf4f2d45af82c56f/VHS_ALG_S1_07_L314_LQ4.gif

jigglypuff314 (jigglypuff314):

solid line shaded to left -> "x <_"

OpenStudy (lovelyanna):

B

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (lovelyanna):

Thanks! :) Last one if its ok?

jigglypuff314 (jigglypuff314):

of course ;)

OpenStudy (lovelyanna):

Thanks! =D Which inequality does this graph represent? A. 3y – 5x < –6 B. 3y – 6x > –6 C. 3y – 6x ≤ –6 D. 3y – 5x ≥ –6 http://static.k12.com/calms_media/media/1312500_1313000/1312943/1/868b5e38cfcfaf4d165e71cc223d2e0043aa1300/HS_A1G2_S1_07_14_114290.jpg

jigglypuff314 (jigglypuff314):

dotted, above -> "y >"

OpenStudy (lovelyanna):

B

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (lovelyanna):

Thanks a bunch!

OpenStudy (lovelyanna):

brb

OpenStudy (lovelyanna):

Ok, I'm back! =D Thanks again! So, how are you today Mad'am? ;P

OpenStudy (lovelyanna):

Or should it be mademoiselle?

jigglypuff314 (jigglypuff314):

lol idk french :P

OpenStudy (lovelyanna):

I don't either :P I just like to address my friends with that for fun ;)

OpenStudy (lovelyanna):

So, how are you?

jigglypuff314 (jigglypuff314):

I'm alright :) I've got a math test coming up that I'm screwed for :P

OpenStudy (lovelyanna):

Oh.. you gonna be ok? Why don't you ask that person that you said he was great at helping you?

jigglypuff314 (jigglypuff314):

It just concepts that I'll have to memorize ;) I'll manage :/

OpenStudy (lovelyanna):

Oh.. I hope so.. that crooked face doesn't convince :P

jigglypuff314 (jigglypuff314):

lol ;) I got a letter to go to an NHS meeting :P

OpenStudy (lovelyanna):

@jigglypuff314 Sorry, my computer disconnected! Congrats! (I don't even know what that means) xD

jigglypuff314 (jigglypuff314):

National Honor Society It's a leadership scholarship thing :P

OpenStudy (lovelyanna):

Oooh.. still don't get it.. :/

OpenStudy (lovelyanna):

brb

OpenStudy (lovelyanna):

Back, sorry I took long @jigglypuff314